What is the diameter of the given circle? (x + 9)2 + (y +13)? = 1296
10 What is the radius of a circle with properties: Center: (-6, 10) Circumference: 1811
Answer #1

Solution. Given equation of circle. Cae+972 +9+13)2 =1296. = (x-(-9)2 + (9-(-131)2 = (36)2. compaining with (2-a)²+(4.b)2=82.